McMahon GFS 2024041606 z Missing - Last was 2024041512 z

McMahon - Last Good GFS was ( 2024041512 z ) and Last Should Be ( 2024041606 z )

Note: this was Checked at 2024-04-16 13:06 UTC.

It is possible that the Data was running Late and may now be available.

Note: this is an automated Post, check data to ensure it is correct!

Kindest Regards,


Sorry, I’ve only just seen that GFS is failing. I’ve been busy with other things today. I’ll try to figure out what’s wrong this evening.

Started to look at this whilst tea is cooking! I use FTP for some tests of whether data is ready and it looks like the FTP access to the GFS data isn’t working. I don’t know if that’s temporary or permanent.

I think I have another way of checking if the data is ready, but that needs a bit of work to get it working so I’ll see how fast I can do that.

The NWS datacenter in College Park has had a major outage due to a chiller failure and lots of over-temp alarms.

They reported (via nco-ops problem tracking)

Additional comments
College Park Datacenter Issue - Update # 4

Key Points

Several storage systems in the College Park data center remain down after Monday afternoon's temperature spike.
Systems have been automatically repairing themselves overnight, but NCO's NetApp Team are working to provide an estimate of when storage systems will be fully recovered.
Additional support was engaged overnight to keep IDP Applications running in Boulder healthy and to investigate mitigating impacts to WPC and OPC operations.
No ETR at this time.
Next update will be at 8am EDT.
Current Known Impacts include:
NCEP Centers' websites hosted in CP are currently inaccessible (including NCO, OPC, EMC, CPC, and NCWCP intranet sites)
WPC, OPC, and CPC's operational product suites are severely degraded
FTPPRD is inaccessible in CP (Customers are able to use as a viable backup in the meantime)
See the NCO IT Status Dashboard for the latest updates and more information

SDM Mary Beth

Its still having issues and even the status page at is having issues. I got the above info by reporting the problem and being added to the mailing list for the incident.

I did a quick search for problems but that didn’t pop up in the results.

I made some good progress at removing FTP before we started to eat. I’m hopeful that I can have it working in about 30 minutes once I’m back at my laptop

The 20240416 12z data is currently downloading using a script that no longer uses FTP. The data should be available to download in about 35 minutes.

I’d actually started the work to stop using FTP some time ago but hadn’t seen it through to completion. Luckily the script fragment I was using to test was easy to fit into the existing download script.